# A Multimodal LLM Approach for Visual Question Answering on Multiparametric 3D Brain MRI

### v2

@@ Abstract
We introduce mpLLM, a prompt-conditioned hierarchical mixture-of-experts (MoE) architecture for visual question answering over multiparametric 3D brain MRI (mpMRI). mpLLM routes across modality-level and token-level projection experts to fuse multiple interrelated 3D modalities, enabling efficient training without image–report pretraining. To address limited image-text paired supervision, mpLLM integrates a synthetic visual question answering (VQA) protocol that generates medically relevant VQA from segmentation annotations, and we collaborate with medical experts for clinical validation. mpLLM outperforms strong medical VLM baselines by 5.3% on average across multiple mpMRI datasets. Our study features three main contributions: (1) the first clinically validated VQA dataset for 3D brain mpMRI, (2) a novel multimodal LLM that handles multiple interrelated 3D modalities, and (3) strong empirical results that demonstrate the medical utility of our methodology. Ablations highlight the importance of modality-level and token-level experts and prompt-conditioned routing.
1 Introduction
Multiparametric MRI (mpMRI) plays a significant role in diagnosing, grading, treating, and assessing treatment responses for brain tumors and other intracranial lesions (Sawlani et al., 2020; Wang et al., 2022a; Cherubini et al., 2016). Describing imaging that involves a complex pattern of brain lesions across multiple regions can be challenging and time-consuming for clinicians. Consequently, several studies have been conducted to develop image recognition and localization models to support clinicians (Ghadimi et al., 2025; Rathore et al., 2018; Wang et al., 2022a; Li et al., 2023c; Osman, 2019).
However, existing models have limited clinical utility because clinicians cannot effectively pose natural language queries about mpMRI. While 3D vision-language models (VLMs) have been developed for other imaging domains, current architectures do not naturally leverage the interdependencies among mpMRI modalities (Li et al., 2023a; Wu et al., 2023; Bai et al., 2024; Xin et al., 2025). Additionally, the standard multi-image approach multiplies the number of vision tokens by the number of images, which significantly increases computational constraints (Wu et al., 2023).
We introduce mpLLM, a prompt-conditioned hierarchical mixture-of-experts (MoE) for VQA over mpMRI. Conditioned on the input question, a router allocates computation across modality-level experts and token-level projection experts to achieve parameter-efficient fusion of multiple interrelated 3D modalities. Unlike modality-specific or modality-agnostic vision encoders, our low-level components are lightweight projection functions that train end-to-end with the language model during fine-tuning. Additionally, the multimodal LLM processes a single fused vision token representation, which dramatically reduces the GPU memory usage.
To address limited image-text paired supervision, we pair mpLLM with a synthetic VQA protocol that derives medically relevant VQA from segmentation annotations, and we obtain clinician validation of both the generated data and model responses. In contrast to prior works, we fine-tune our model using next-token prediction directly on the VQA dataset without pretraining on a paired imaging-report dataset. We also train a multi-task head end-to-end with the multimodal LLM for improved task proficiency and more reliable evaluation. In summary, our research makes these key contributions:
  1. 1.
    In collaboration with medical experts, we introduce a synthetic VQA protocol that produces the first clinically validated VQA dataset for 3D brain mpMRI.
  2. 2.
    We design mpLLM, a multimodal LLM that uses a prompt-conditioned hierarchical MoE to effectively leverage the interdependence between 3D modalities in mpMRI.
  3. 3.
    Strong empirical results that support our methodology as a foundation for future research with multimodal LLMs in brain mpMRI.
@@ 1 Introduction
Figure 1: High-level comparison between LLaVA-Med and mpLLM. While LLaVA-Med uses a standard projection layer, our method uses a hierarchical MoE block which ingests both the prompt and imaging to produce prompt-conditioned vision tokens that leverage all the 3D modalities.
2 Related Work
@@
Although several 3D VLMs exist for natural images (Zhu et al., 2024b; Li et al., 2024b; Zhu et al., 2023), they require access to extremely large annotated datasets, which are often unavailable in medical contexts. While a few 3D VLMs have been developed for medical imaging, these methods have certain limitations. In one recent paper, researchers adapted the LLaVA-Med architecture to utilize spatial pooling and pretrain a 3D vision encoder with 700k radiology images (Bai et al., 2024). In another paper, researchers pretrain segmentation modules to generate brain imaging reports (Lei et al., 2024). In recent work, researchers exploit vision-language pretraining for CT report generation (Liu et al., 2023a; Chen & Hong, 2024; Blankemeier et al., 2024; Xin et al., 2025; Cao et al., 2025). However, these prior works assume a large paired imaging-report pretraining dataset, which is infeasible to collect and imposes a significant training burden.
@@
Previous work in MoE has concentrated on training and inference efficiency (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Zoph et al., 2022; Liu et al., 2024a), transfer learning (Li et al., 2022; Zhong et al., 2022), class imbalance (Han et al., 2024), and multi-domain information (Zhang et al., 2024). There have also been earlier efforts with multimodal LLMs, covering sparsity learning (Lin et al., 2024), task interference (Shen et al., 2025), and embedding models (Li & Zhou, 2024). Related to our work, several studies have employed MoE with VLMs to select between vision encoders and vision-language projections (Li et al., 2025; Zong et al., 2024; Wang et al., 2023; Ma et al., 2025). However, these studies address two modalities and do not account for interactions between different 3D image modalities, which present additional challenges our work seeks to address.
MoE also has various applications in the medical field. These applications include addressing missing modalities (Yun et al., 2025; Novosad et al., 2024; Liu et al., 2024d), fairness (Wang et al., 2025), pediatric care (Huy et al., 2025), parameter reduction and efficiency (Jiang et al., 2024; Nathani et al., 2024), and super resolution (Lin et al., 2021). Additionally, several studies have focused on the segmentation of multimodal medical imaging (Zhang et al., 2025a; Jiang & Shen, 2024). However, no existing research has explored using MoEs for multiple interrelated 3D image modalities. This area is particularly complex due to the need to project multiple interrelated vision modalities into the language modality.
@@
One of the primary challenges in report generation is evaluation: lexical metrics such as BLEU, ROUGE‑L, and BERTScore have been shown to correlate poorly with radiologist evaluations (Yu et al., 2023). In contrast, VQA allows for more granular and interpretable model evaluation. While there are several medical VQA datasets, many focus on 2D imaging (Liu et al., 2024b; 2021; He et al., 2020; Lau et al., 2018). In a prior work, researchers used a scene graph generator to generate surgical VQA (Yuan et al., 2024). In a recent work, researchers extracted multi-task questions from structured lung cancer screening data (Niu et al., 2025). However, there is no existing VQA dataset for 3D brain mpMRI due to a significant lack of source data for VQA extraction. In our work, we remedy this by leveraging publicly available segmentation annotations as source data.
3 Methodology
Figure 2: Detailed overview of our mpLLM pipeline.
Brain mpMRI has several 3D imaging modalities. Given a specific 3D image modality $I_{m}\in R^{C\times D\times H\times W}$, where $m$, $C$, $D$, $H$, and $W$ represent the modality, channel, depth, height, and width respectively, we derive the image modality embedding $v_{m}=h(I)\in R^{N_{I}\times d_{I}}$. Here, $h$ denotes the vision encoder, $N_{I}$ indicates the number of image tokens produced by the vision encoder, and $d_{I}$ represents the vision encoder token embedding dimension. We pass the image modality embeddings through a spatial pooling layer, which reduces the number of tokens, and concatenate them for all the image modalities before passing them to the hierarchical MoE.
Given the concatenated image modality embeddings $v\in R^{N_{I}\times N_{m}\times d_{I}}$ where $N_{m}$ represents the number of image modalities, we derive the projected image modality embedding $e=MoE(v,t)\in R^{N_{I}\times d_{T}}$ where $MoE$ denotes the hierarchical MoE, $t$ represents the text prompt, and $d_{T}$ represents the LLM embedding layer dimension. $(t,e)$ is then provided as a soft prompt to the LLM for multi-task prediction and text generation. A detailed visualization of our approach can be seen in Figure 2.
3.1 Hierarchical mixture-of-experts for multiparametric MRI projection
@@
Our hierarchical MoE architecture includes a high-level router $h$ that assigns weights over a set of high-level experts $\{\mathcal{E}^{(h)}_{1},...,\mathcal{E}^{(h)}_{N}\}$, where $N$ is the number of high-level experts. These experts operate at the image modality and image token levels. The router is implemented as a two-layer MLP that takes as input the final hidden state of the language model corresponding to the text prompt $t$. It produces a normalized weight distribution over experts: $\pi^{(h)}(t)=softmax(h(t))\in R^{N}$. Since task information is embedded within the text prompt, the router implicitly infers the task, enabling the high-level experts to specialize in different task proficiencies.
@@
Our hierarchical MoE includes high-level experts operating at different granularity levels: image modality-level and image token-level. Each high-level expert consists of a two-layer MLP low-level router $l$ and a set of associated low-level experts $\{W^{(l)}_{1},...,W^{(l)}_{M}\}$ where $M$ is the number of low-level experts.
The image modality-level expert takes as input the concatenated [CLS] tokens from all image modalities (e.g., T1, T2), and outputs image modality-level weights over the corresponding low-level experts: $\pi^{(l)}(v)=sigmoid(l(v))\in R^{M}$. In contrast, the token-level expert receives the $i$-th position image tokens from all image modalities as input and outputs image token-level weights over the low-level experts for position $i$: $\pi^{(l)}(v)=sigmoid(l(v))\in R^{M\times N_{I}}$. As discussed in prior work (Li et al., 2024a), providing weights at different granularities improves task performance by enhancing domain generalizability.
@@
Each image modality is passed through both low-level experts and then summed embedding dimension-wise. To balance contributions from each image modality and prevent mode collapse towards a single image modality, the low-level router generates weights that total 1 for both experts. The overall formulation for the hierarchical MoE is as follows:
$MoE(v,t)=\sum_{n\in[N]}\pi^{(h)}_{n}\sum_{m\in[N_{m}]}\pi^{(l,n)}_{m}W_{m,n}(v_{m})+(1-\pi^{(l,n)}_{m})W_{shared,n}(v_{m})$ (1)
where $n$ represents the high-level expert and $m$ represents the image modality. We considered several different hyperparameters with our MoE implementation, including the number of high-level experts, prompt information in high-level and/or low-level routers, the number of layers in the low-level experts, and others. Our validation experiments on the GLI dataset found that the optimal number of high-level experts was 16, corresponding to the number of labels times the number of tasks.
@@
Table 1: Statistics for the synthetic VQA datasets.
3.2 Training objectives
3.2.1 Synthetic VQA protocol
Because of the lack of brain mpMRI VQA data, we propose a novel method of synthetic VQA generation that leverages the publicly available brain mpMRI segmentation data. To generate relevant VQA data, we consult with clinicians to identify important topics that can be extracted from the label masks, focusing on mask volume relative to brain volume (Kaifi, 2023), brain region localization (Lau et al., 2018), shape (Ismail et al., 2018), and spread (Islam et al., 2019). For each label mask, we compute the quantities using standard formulas and validate the thresholds with synthetic masks and a subset of data. To emulate the subjectivity found in medical reports, we categorize each of the quantities based on their magnitude using terminology similar to that found in medical reports. Rather than using an LLM, we employ a rules-based method to assign medical terms to the quantities, ensuring our approach is clinically relevant and highly reliable. We assign “N/A” if the label is not found.

### v3

# Multimodal LLM With Hierarchical Mixture-of-Experts for VQA on 3D Brain MRI

@@ Abstract.
Multiparametric 3D brain MRI (mpMRI) is central to neuroradiology, but producing tumor location, appearance, size, and involvement of critical structures for neurosurgical planning remains challenging. We introduce mpLLM, a multimodal LLM for visual question answering (VQA) on mpMRI that produces clinically interpretable tumor descriptors (e.g., volume, morphology, extent, and coarse localization) as an adjunct to clinical expertise for referring neurosurgeons. mpLLM uses a prompt-conditioned hierarchical mixture-of-experts (MoE) to fuse multiple 3D sequences via routing over modality- and token-level projection experts, enabling data-efficient end-to-end training without large-scale image–report pretraining. To address limited paired image–text supervision, we propose a synthetic VQA protocol that derives clinically grounded questions and answers from expert segmentation annotations and is validated with radiologist collaboration. Across multiple mpMRI datasets, mpLLM improves over strong medical VLM baselines by +5.5 points on average (+9.1% relative) and increases radiologist-rated clinical acceptability by +15.9 points (+46.6% relative). Our study features three main contributions: (1) the first VQA dataset for 3D brain mpMRI, (2) a hierarchical MoE architecture for joint reasoning over interrelated 3D sequences, and (3) expert-supported evidence of clinical utility. Source code is available at https://github.com/arvindmvepa/mpllm, and we will release the dataset upon publication.
Multimodal LLM, Mixture-of-Experts, VQA, Synthetic Data, Medical Imaging
copyright: noneccs: Applied computing Health informaticsccs: Computing methodologies Neural networksccs: Computing methodologies Natural language generationccs: Computing methodologies Computer vision
@@
Figure 1. High-level comparison between LLaVA-Med and mpLLM. While LLaVA-Med uses a standard projection layer, our method uses a hierarchical MoE block which ingests both the prompt and imaging to produce prompt-conditioned vision tokens that leverage all the 3D modalities.
High-level comparison between LLaVA-Med and mpLLM. While LLaVA-Med uses a standard projection layer, our method uses a hierarchical MoE block which ingests both the prompt and imaging to produce prompt-conditioned vision tokens that leverage all the 3D modalities.
1. Introduction
Multiparametric MRI (mpMRI) plays a significant role in diagnosing, grading, treating, and assessing treatment responses for brain tumors and other intracranial lesions (Sawlani et al., 2020; Wang et al., 2022a; Cherubini et al., 2016). Producing clear, consistent descriptions of tumor size, morphology, extent, and localization is time-consuming, yet critical for downstream decision-making. As a result, a growing body of work has developed recognition and localization models to support clinical interpretation (Ghadimi et al., 2025; Rathore et al., 2018; Wang et al., 2022a; Li et al., 2023c; Osman, 2019).
A key unmet need is interactive, queryable assistance: neurosurgical planning and counseling depend not only on detecting tumors but on communicating clinically interpretable descriptors—where the lesion is, what it looks like, how large it is, and their involvement of critical structures. However, existing models have limited clinical utility because clinicians cannot effectively pose natural language queries about mpMRI. While 3D vision-language models (VLMs) have been developed for other imaging domains, current architectures do not naturally leverage the interdependencies among mpMRI modalities (Li et al., 2023a; Wu et al., 2023; Bai et al., 2024; Xin et al., 2025). Additionally, the standard multi-image approach multiplies the number of vision tokens by the number of images, which significantly increases computational constraints (Wu et al., 2023).
We introduce mpLLM, a prompt-conditioned hierarchical mixture-of-experts (MoE) for VQA over mpMRI. Our approach is an extension of the LLaVa architecture (Liu et al., 2023b). mpLLM replaces the standard single projection with a prompt-conditioned MoE projection that routes across modality-level and token-level projection experts to fuse multiple 3D sequences into an efficient representation for the language model. Unlike approaches that require training separate modality-specific encoders, we use lightweight projection functions that train end-to-end with the language model during fine-tuning.
To address limited paired image–text supervision in neuroimaging, we pair mpLLM with a synthetic VQA protocol that derives clinically grounded question–answer pairs from expert segmentation annotations, developed in collaboration with radiologists to ensure clinical relevance. In contrast to prior work, we fine-tune directly on the resulting VQA supervision via next-token prediction without pretraining on paired image–report data, and we train an end-to-end multi-task head to improve task proficiency and support reliable evaluation. In summary, our research makes these key contributions:
  1. (1)
    In collaboration with radiologists, we introduce a synthetic VQA protocol that produces the first VQA dataset for 3D brain mpMRI.
  2. (2)
    We design mpLLM, a multimodal LLM with prompt-conditioned hierarchical MoE routing that leverages interdependencies among 3D mpMRI modalities.
  3. (3)
    We provide empirical evidence of clinical utility through expert-supported validation and clinically relevant downstream evaluation.
Figure 2. Detailed overview of our mpLLM pipeline.
Detailed overview of our mpLLM pipeline.
2. Related Work
@@
Although several 3D VLMs exist for natural images (Zhu et al., 2024b; Li et al., 2024b; Zhu et al., 2023), they require access to extremely large annotated datasets, which are often unavailable in medical contexts. While a few 3D VLMs have been developed for medical imaging, these methods have certain limitations. In one recent paper, researchers adapted the LLaVA-Med architecture to utilize spatial pooling and pretrain a 3D vision encoder with 700k radiology images (Bai et al., 2024). In another paper, researchers pretrain segmentation modules to generate brain imaging reports (Lei et al., 2024). In recent work, researchers exploit vision-language pretraining for CT report generation (Liu et al., 2023a; Chen and Hong, 2024; Blankemeier et al., 2024; Xin et al., 2025; Cao et al., 2025; Li et al., 2025a). However, these prior works assume a large paired imaging-report pretraining dataset, which is infeasible to collect and imposes a significant training burden.
@@
Previous work in MoE has concentrated on training and inference efficiency (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Zoph et al., 2022; Liu et al., 2024a), transfer learning (Li et al., 2022; Zhong et al., 2022), class imbalance (Han et al., 2024), and multi-domain information (Zhang et al., 2024). There have also been earlier efforts with multimodal LLMs, covering sparsity learning (Lin et al., 2024), task interference (Shen et al., 2025), and embedding models (Li and Zhou, 2024). MoE also has various applications in the medical field. These applications include addressing missing modalities (Yun et al., 2025; Novosad et al., 2024; Liu et al., 2024d), fairness (Wang et al., 2025), pediatric care (Huy et al., 2025), parameter reduction and efficiency (Jiang et al., 2024; Nathani et al., 2024), super resolution (Lin et al., 2021), and segmentation of multimodal medical imaging (Zhang et al., 2025a; Jiang and Shen, 2024).
Related to our work, several studies have employed MoE with VLMs to select between vision encoders and vision-language projections (Li et al., 2025b; Zong et al., 2024; Wang et al., 2023; Ma et al., 2025). In CuMo (Li et al., 2025b), the authors incorporate MoE in the MLP layers in the vision encoder as well as projection function; however, they do not utilize prompt-conditioning, do not focus on combining multiple inputs, and require a complex training scheme (our approach requires a single stage of instruction fine-tuning). In MoVA (Zong et al., 2024), the authors utilize MoE with task-specific vision encoders. In practice, training vision encoders is challenging and resource intensive in the medical space (Rui et al., 2024). Improving the LLaVa projection function with a MoE block uses far less parameters and can be trained end-to-end with the LLM using next-token prediction. The previous approach also does not focus on combining multiple inputs. Finally, in 3D-MoE (Ma et al., 2025), the authors utilize MoE within the LLM and do not focus on combining multiple inputs. Additionally, no existing research has explored using MoE to project multiple interrelated 3D image modalities into the language embedding space.
@@
One of the primary challenges in report generation is evaluation: lexical metrics such as BLEU, ROUGE‑L, and BERTScore have been shown to correlate poorly with radiologist evaluations (Yu et al., 2023). In contrast, VQA allows for more granular and interpretable model evaluation. While there are several medical VQA datasets, many focus on 2D imaging (Liu et al., 2024b, 2021; He et al., 2020; Lau et al., 2018). In a prior work, researchers used a scene graph generator to generate surgical VQA (Yuan et al., 2024). In a recent work, researchers extracted multi-task questions from structured lung cancer screening data (Niu et al., 2025). While there are existing VQA datasets for 2D brain MRI (Safari et al., 2025; Shehzad et al., 2025; Peng et al., 2025), there are no existing VQA datasets for 3D brain mpMRI due to a significant lack of source data for VQA extraction. In our work, we remedy this by collaborating with radiologists to generate clinically relevant synthetic data from expert annotated segmentation annotations and leverage ChatGPT-4o for dataset augmentation. Other prior works have also leveraged commercial LLMs to generate VQA data from expert annotations and study metadata in the medical setting (Zhang et al., 2023b; Gautam et al., 2024; Butsanets et al., 2025). To contextualize dataset reliability and demonstrate practical value, we report data quality metrics, radiologist acceptability ratings of model-generated responses, and performance on clinically relevant downstream tasks.
3. Methodology
Brain mpMRI has several 3D imaging modalities. For a given modality $m$, let $I_{m}\in\mathbb{R}^{C\times D\times H\times W}$ denote the corresponding 3D volume, where $C$, $D$, $H$, and $W$ represent the channel, depth, height, and width, respectively. A 3D vision encoder $h_{\text{vis}}$ maps each modality to a sequence of image tokens $v_{m}=h_{\text{vis}}(I_{m})\in\mathbb{R}^{T\times d_{V}}$, where $T$ is the number of image tokens and $d_{V}$ is the vision embedding dimension. We apply spatial pooling to reduce the token length (reusing the symbol $T$ for the pooled length for simplicity) and then concatenate the pooled tokens across the $M$ image modalities before passing them to the hierarchical MoE.
Let $v\in\mathbb{R}^{M\times T\times d_{V}}$ denote the concatenated image modality embeddings. The hierarchical MoE projects these embeddings into the LLM space, $e=\mathrm{MoE}(v,t)\in\mathbb{R}^{M\times T\times d_{T}}$, where $\mathrm{MoE}$ denotes the hierarchical MoE block, $t$ represents the text prompt, and $d_{T}$ is the LLM embedding dimension. In practice we flatten the modality and token dimensions and provide $(t,e)$ as a soft prompt to the LLM for multi-task prediction and text generation. A detailed visualization of our approach can be seen in Figure 2.
3.1. Hierarchical mixture-of-experts for multiparametric MRI projection
In what follows, we use the term expert exclusively for projection modules that map vision features into the LLM embedding space, and the term router for the MLPs that output mixing weights over experts. Each high-level expert block therefore consists of a router together with its associated projection experts.
@@ 3.1. Hierarchical mixture-of-experts for multiparametric MRI projection
Our hierarchical MoE architecture includes a high-level router $r^{(h)}$ that assigns weights over a set of high-level expert blocks $\{\mathcal{E}^{(h)}_{1},\dots,\mathcal{E}^{(h)}_{H}\}$, where $H$ is the number of high-level experts. These expert blocks operate at the image modality and image token levels. The router is implemented as a two-layer MLP that takes as input the final hidden state of the language model corresponding to the text prompt $t$. It produces a normalized weight distribution over expert blocks: $\pi^{(h)}(t)=\mathrm{softmax}(r^{(h)}(t))\in\mathbb{R}^{H}$. Since task information is embedded within the text prompt, the router implicitly infers the task, enabling high-level experts to specialize in different task proficiencies.
@@
Our hierarchical MoE includes high-level experts operating at different granularity levels: image modality-level and image token-level. Each high-level expert block consists of a two-layer MLP low-level router $r^{(l)}$ and an associated set of low-level projection experts $\{W^{(l)}_{1},\dots,W^{(l)}_{L}\}$, where $L$ is the number of low-level experts within the block.
The image modality-level expert block uses a low-level router that takes as input the concatenated [CLS] tokens from all image modalities (e.g., T1, T2) and outputs modality-level weights over the corresponding low-level experts: $\pi^{(l)}_{\text{mod}}(v)=\sigma(r^{(l)}_{\text{mod}}(v))\in\mathbb{R}^{L}$. In contrast, the token-level expert block uses a low-level router that receives, for each token position $i$, the $i$-th image tokens from all modalities and outputs token-level weights $\pi^{(l)}_{\text{tok}}(v)=\sigma(r^{(l)}_{\text{tok}}(v))\in\mathbb{R}^{L\times T}$. As discussed in prior work (Li et al., 2024a), providing weights at different granularities improves task performance by enhancing domain generalizability.
@@
Each image modality is passed through both low-level experts and then summed embedding dimension-wise. The overall formulation for the hierarchical MoE is as follows:
(1) $\mathrm{MoE}(v,t)$ $=\sum_{h=1}^{H}\pi^{(h)}_{h}(t)\,\cdot$
$\quad\sum_{m=1}^{M}\Big(\alpha^{(h)}_{m}\,W^{(h)}_{m}(v_{m})+\beta^{(h)}_{m}\,W^{(h)}_{\text{shared}}(v_{m})\Big)$
where $h$ indexes the high-level expert blocks, $m$ indexes the image modalities, and $\alpha^{(h)}_{m}$ and $\beta^{(h)}_{m}$ are the modality-specific and modality-agnostic weights produced by the corresponding low-level router (which sum to one within an expert block $h$). Our validation experiments on the GLI dataset found that the optimal number of high-level experts was 16, corresponding to the number of labels times the number of tasks.
@@
Table 1. Comparison of task performance for all models on all datasets with accuracy metric with standard deviation.
@@
[…]
3.2. Training objectives
3.2.1. Synthetic VQA protocol
Because of the lack of brain mpMRI VQA data, we propose a novel method of synthetic VQA generation that leverages the expert annotated brain mpMRI segmentation data. We perform the following steps: 1) compute continuous quantities from masks, 2) map to quantities to report-style discrete labels, and 3) insert discrete labels into paraphrased QA templates.
To generate relevant VQA data, we consult with radiologists to identify important topics that can be extracted from the label masks, focusing on mask volume relative to brain volume (Kaifi, 2023), brain region localization (Lau et al., 2018), shape (Ismail et al., 2018), and spread (Islam et al., 2019). For each label mask, we compute the quantities using standard formulas and validate the thresholds with synthetic masks and a subset of data. To emulate the subjectivity found in medical reports, we categorize each of the quantities based on their magnitude using terminology similar to that found in medical reports. We employ a rules-based method to assign medical terms to the quantities. We assign “N/A” if the label is not found.